Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2025 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 79     Entry time: 02 Jun 2025, 07:59
Author: Yoonhee Kim 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 9.3 keV, 1600-1700 uJ photon flux, beam down once due to Big Brother intervention. But the signal on HIREX was less than 10 counts.

Optical laser delivery

  • ns laser for particle illumination inside the chamber
  • another ns laser to visualize the tip of capillary on top of the chamber

Achievements/Observations

  • Measured granulovirus

  • Hit rate of defocused beam was similar or less than with focused beam, Q range is smaller than with focused beam.

  • Without dryer, hit rate is more than twice.

Issues

  • BIG BROTHER is intervened due to HIREX and beam was down due to that. But the count on HIREX with Al1.2 was less than 100 count.

Instrument status

  • AGIPD HV off.
  • AGIPD gate valve closed and pinned.
  • Attenuators set to thickest diamond.
  • Vacuum valves closed.
  • AGIPD breaks enguaged
  • ns Lasers off
  • HIREX is out
ELOG V3.1.4-7c3fd00